### v2

# A Physics-Informed Spatiotemporal Deep Learning Framework for Turbulent Systems

Abstract
Fluid thermodynamics underpins atmospheric dynamics, climate science, industrial applications, and energy systems. However, direct numerical simulations (DNS) of such systems can be computationally prohibitive. To address this, we present a novel physics-informed spatiotemporal surrogate model for Rayleigh-Bénard convection (RBC), a canonical example of convective fluid flow. Our approach combines convolutional neural networks, for spatial dimension reduction, with an innovative recurrent architecture, inspired by large language models, to model long-range temporal dynamics. Inference is penalized with respect to the governing partial differential equations to ensure physical interpretability. Since RBC exhibits turbulent behavior, we quantify uncertainty using a conformal prediction framework. This model replicates key physical features of RBC dynamics while significantly reducing computational cost, offering a scalable alternative to DNS for long-term simulations.
Keywords: Long-range forecasting, state-space modeling, physics-informed inference, fluid thermodynamics
1 Introduction & Background
Modeling fluid thermodynamics is critical for advancing our understanding of complex physical systems, including atmospheric dynamics (Trenberth & Stepaniak 2003), climate science (Held & Suarez 1994), and industrial processes (Singh & Heldman 2014). Accurate thermodynamic models enable the prediction of phase transitions, heat transfer, and fluid motion, key to optimizing engineering systems and mitigating environmental impacts (Gyftopoulos & Beretta 2012). However, Direct Numerical Simulation (DNS) is computationally expensive due to the wide range of spatial and temporal scales and strong nonlinearities involved (Xu 2015). As a result, statistical surrogate models have emerged as promising alternatives, capable of learning from limited data and enabling longer simulations at significantly lower cost.
Surrogates are effective for modeling turbulence only if they can capture the nonlinear, multiscale structure of the governing dynamics. To this end, reduced-order models such as Proper Orthogonal Decomposition and Koopman operator theory have provided principled, interpretable, and computationally efficient frameworks for simulating dominant flow structures (Rowley 2005, Brunton et al. 2021). More recently, operator learning approaches—such as the Fourier Neural Operator—have shown promise in approximating Partial Differential Equations (PDE) solution operators, scaling to high-resolution turbulent flows (Li et al. 2020). These methods aim to learn solution operators rather than simple input-output maps. However, most are deterministic and lack mechanisms for uncertainty quantification, limiting their reliability in sensitive or chaotic regimes.
An alternative paradigm uses supervised learning to emulate turbulent flows directly from data. Convolutional Neural Networks (CNNs) have been used to predict flow fields from boundary conditions or coarse inputs, while Recurrent Neural Networks (RNNs) model temporal evolution (Guo et al. 2016, Maulik et al. 2021). Convolutional Recurrent Neural Networks (CRNNs) — which combine CNN-based spatial encoding with RNN-based temporal modeling — have shown success in spatiotemporal forecasting of turbulent flows. These models typically employ an encoder to compress spatial structure, an RNN to model dynamics in reduced space, and a decoder to reconstruct high-dimensional output. Applications include two-dimensional cylinder flows (Beiki & Kamali 2023), turbulence reconstruction from sparse data (Fukami & Fukagata 2021), and short-term forecasting in low-turbulence settings (Mohan et al. 2019, Pandey et al. 2022, Straat et al. 2025). These models are however currently limited in that their architectures are suboptimal in long-range forecasting and are not formulated as statistical models, so they don’t allow formal uncertainty quantification.
In parallel, a growing body of work in scientific machine learning seeks to improve model interpretability by embedding physical laws directly into learning algorithms. Physics-Informed Neural Networks (PINNs) incorporate PDE constraints into inference via penalized optimization, leading to predictions that better align with physical theory (Cuomo et al. 2022, Raissi et al. 2019). While PINNs have been applied to systems like the 1D Burgers’ equation, laminar flows, and Reynolds-averaged Navier-Stokes, their extension to fully turbulent, high-dimensional systems remains an open challenge (Rao et al. 2020, Thuerey et al. 2020). Key limitations include a focus on weakly turbulent regimes, limited long-range accuracy, and lack of the quantification of predictive uncertainty.
In this work, we propose a novel Physics-Informed Convolutional Recurrent Neural Network (PI-CRNN) for efficient, long-range forecasting of fluid thermodynamic systems. Our model integrates four key components: (1) spatial feature extraction via a convolutional autoencoder; (2) temporal modeling through a sequence-to-sequence ConvLSTM architecture; (3) physics-informed inference by penalizing violations of mass, momentum, and energy conservation; and (4) uncertainty quantification using conformal prediction (Shi et al. 2015, Bonas et al. 2024). While previous work has explored CRNNs and PINNs separately, to our knowledge, this is the first model that embeds a ConvLSTM within the reduced space of a convolutional autoencoder while enforcing PDE constraints and providing distribution-free prediction intervals. PI-CRNN model is intended to be used as an offline-trained statistical surrogate, which is crucial for a variety of scientific and engineering applications where long time series of a turbulent physical process are needed but DNS is considerably more expensive and time-consuming.
Our testbed is Rayleigh-Bénard convection (RBC), a canonical example in fluid thermodynamics featuring nonlinear heat-driven flow between differentially heated surfaces (Rayleigh 1916). RBC exhibits high-dimensional turbulence and is highly sensitive to initial conditions, making it ideal for evaluating long-term forecast accuracy and uncertainty quantification. While the governing physical equations, i.e., the data-generating process, are deterministic, this sensitivity results in de facto stochastic behavior. We incorporate physical constraints by enforcing the Navier-Stokes equations via penalized optimization (Navier 1822, Stokes 1845), and quantify uncertainty through conformal prediction, which guarantees calibrated prediction intervals under minimal assumptions. The turbulent nature of RBC, coupled with the computational cost of DNS, motivates the development of a statistical surrogate capable of capturing the spatiotemporal dynamics and associated uncertainties at a fraction of the computational expense. PI-CRNN offers a unified, principled framework for surrogate modeling of complex thermodynamic systems, balancing computational efficiency, physical plausibility, and statistical rigor.
The remainder of the manuscript is structured as follows: Section 2 summarizes the turbulent convection data and governing equations; Section 3 introduces the PI-CRNN model; Section 4 presents the physics-informed learning framework; Section 5 discusses uncertainty quantification; Section 6 reports the results; and Section 7 concludes with a discussion of limitations and future directions.
2 Data
This work utilizes simulated data from a two-dimensional RBC flow for dry air, a gravity-driven turbulent flow between two horizontal plates exhibiting a temperature gradient. The dataset, generated via DNS, includes horizontal and vertical velocity components, pressure, and temperature with high spatial and temporal resolution. RBC is governed by the Navier-Stokes equations under the Boussinesq approximation, which encode mass conservation, momentum, and energy conservation for a thermally stratified fluid (Navier 1822, Boussinesq 1903, Rayleigh 1916). The Rayleigh number for this data (a quantity identifying the amount of turbulence in the system) is $\text{Ra = 2.54}\times 10^{8}$. A low Ra (less than $10^{6}$) corresponds to laminar (smooth) flow, while $\text{Ra}=2.54\times 10^{8}$ indicates strongly turbulent convection, making the dataset both scientifically realistic and statistically challenging. This data offers an ideal testbed: the underlying physics are well established, yet the resulting flow is high-dimensional, nonlinear, and turbulent, posing a significant challenge for long-range forecasting. This motivates the development of a surrogate model that can capture both uncertainty and dynamics efficiently. Further technical details on the data and governing equations are in Section S.2 of the supplementary material.
The data was generated on a $n_{x}\times n_{z}$ grid, $n_{x}=n_{z}=256$, with a temporal resolution of $\Delta_{t}=0.1$ seconds for a total of T=2,500T=2{,}500 time steps. We denote the RBC spatiotemporal process with $d=4$ covariates observed on a $n_{x}\times n_{z}$ grid by $\bm{Y}_{t}(x,z)=\left(u_{t}(x,z),w_{t}(x,z),p_{t}(x,z),\theta_{t}(x,z)\right)$, where $u_{t}(x,z)$ and $w_{t}(x,z)$ are the horizontal and vertical velocity components, respectively, $p_{t}(x,z)$ is pressure, $\theta_{t}(x,z)$ is temperature, $(x,z)$ are spatial coordinates, and $t=1,\ldots,T$. For simplicity, we hereto omit the spatial coordinates, so we denote one observation over the entire spatial domain by $\bm{Y}_{t}=\left(u_{t},w_{t},p_{t},\theta_{t}\right)\in\mathbb{R}^{n_{x}\times n_{z}\times d}$. Since the temporal modeling approach we propose considers sequences of data, e.g., $\left(\bm{Y}_{3},\bm{Y}_{4}\right)=f\left(\bm{Y}_{1},\bm{Y}_{2}\right)$”, we denote the sequence of $b$ observations before $t$ by $\mathcal{Y}^{-}_{t;b}=\left(\bm{Y}_{t-b},\ldots,\bm{Y}_{t-2},\bm{Y}_{t-1}\right)$ and the sequence of $a$ observations starting at $t$ by $\mathcal{Y}^{+}_{t;a}=\left(\bm{Y}_{t},\bm{Y}_{t+1},\ldots,\bm{Y}_{t+a-1}\right)$.
3 Methods
As detailed in Section 2, $\bm{Y}_{t}$ is a high-dimensional process, where one observation consists of $D=n_{x}\times n_{z}\times d$ points. At $\text{Ra}=2.54\times 10^{8}$, RBC exhibits turbulent dynamics, making direct modeling of $\bm{Y}_{t}$ both computationally expensive and statistically challenging. To address this, we model the temporal evolution of $\bm{Y}_{t}$ through a reduced space representation, which we denote by $\bm{\tilde{Y}}_{t}$, of dimensions $\tilde{D}\ll D$. Indeed, $\bm{\tilde{Y}}_{t}$ is intended to capture the essential spatial structure of $\bm{Y}_{t}$ while enabling efficient forecasting. This reduced representation naturally leads to a state-space formulation, where spatial compression and temporal evolution can be modeled separately.
The state-space model is, formally,
Observation Equation: $\quad\bm{Y}_{t}=\mathcal{SD}\left(\underbrace{\mathcal{SE}\left(\bm{Y}_{t}\mid\bm{W}_{\text{en}}\right)}_{\bm{\tilde{Y}}_{t}}\mid\bm{W}_{\text{de}}\right)+\bm{\eta}_{t},$ (1a)
State Equation: $\quad\bm{\tilde{Y}}_{t}=\mathcal{T}\left(\bm{\tilde{Y}}_{t-1},\ldots,\bm{\tilde{Y}}_{t-b}\mid\bm{W}^{\text{time}}\right)+\tilde{\bm{\epsilon}}_{t},$ (1b)
where $\bm{\tilde{Y}}_{t}\in\mathbb{R}^{\tilde{n}_{x}\times\tilde{n}_{z}\times\tilde{d}}$, $\tilde{D}=\tilde{n}_{x}\times\tilde{n}_{z}\times\tilde{d}$, is the reduced space representation of $\bm{Y}_{t}$. In (1a), $\mathcal{SE}(\cdot)$ denotes the spatial encoder and $\mathcal{SD}(\cdot)$ denotes the spatial decoder, so the composition of the two is an autoencoder as a (possibly nonlinear) function of some parameters $\bm{W}^{\text{space}}$. Lastly, $\bm{\eta}_{t}$ is the error term associated with the data reduction which we assume to be Gaussian with mean zero. In (1b), $\mathcal{T}(\cdot)$ models the temporal evolution of $\bm{\tilde{Y}}_{t}$ as a (possibly nonlinear) function of some parameters $\bm{W}^{\text{time}}$. The error term $\tilde{\bm{\epsilon}}_{t}$ captures the prediction uncertainty in reduced space and we assume it to be Gaussian with mean zero. Diagnostics on the Gaussian assumptions are available in Section S.6 of the supplementary material. This modeling approach allows for 1) the complexity of the temporal structure of $\bm{Y}_{t}$ to be greatly simplified in reduced space, since $\bm{\tilde{Y}}_{t}$ only retains key features and 2) the computational burden to be greatly reduced, since the dimensions of $\tilde{\bm{Y}}_{t}$ are $\frac{\tilde{D}}{D}$% of those of $\bm{Y}_{t}$.
In the observation equation (1a), we implement a convolutional autoencoder (CAE), a spatially-aware neural network model where a spatial encoder compresses the input $\bm{Y}_{t}$ into $\bm{\tilde{Y}}_{t}$ (reduced space), while a spatial decoder inputs $\bm{\tilde{Y}}_{t}$ and returns it to the original dimensions $D$ (Masci et al. 2011). As such, if $\bm{Y}_{t}\approx\mathcal{SD}(\bm{\tilde{Y}}_{t}\mid\bm{W}_{\text{de}})$ then the spatial encoder $\mathcal{SE}(\cdot)$ successfully compressed the features of $\bm{Y}_{t}$ into $\bm{\tilde{Y}}_{t}$. In the state equation (1b), the temporal structure of $\bm{\tilde{Y}}_{t}$ is modeled in time with a RNN $\mathcal{T}(\cdot)$ using a sequence-to-sequence framework, i.e., a model that inputs a sequence and outputs a sequence (not necessarily of the same length). We denote the parameters in the temporal model by $\bm{W}^{\text{time}}$ and all the parameters in our spatiotemporal model (1) by $\bm{W}=(\bm{W}^{\text{space}},\bm{W}^{\text{time}})$.
In the following Sections 3.1 and 3.2 we detail the CAE and RNN, respectively. Section 3.3 summarizes the spatiotemporal model.
3.1 Spatial Model: Convolutional Autoencoder
Throughout this section we omit the $t$ subscript from $\bm{Y}_{t}$ for simplicity. We model the spatial structure of $\bm{Y}$ with a CAE, whose key principle is to first project the input onto a reduced space, $\tilde{\bm{Y}}\in\mathbb{R}^{\tilde{n}_{x}\times\tilde{n}_{z}\times\tilde{d}}$, referred to as the spatial encoder, and subsequently projected back onto $\mathbb{R}^{n_{x}\times n_{z}\times d}$, referred to as the spatial decoder. Figure 1 provides an illustration for the CAE model’s architecture, where yellow blocks represent the encoder and blue block represent the decoder. Both the spatial encoder and spatial decoder are nonlinear functions and the dimension-reconstruction in the decoder is not generally the inverse of the dimension-reduction performed in the encoder. In CAEs the input undergoes convolutions, followed by a nonlinear activation function $g$ such as the rectified linear unit, sigmoid, or hyperbolic tangent function (Goodfellow et al. 2016).
The spatial encoder consists of $\ell=1,\ldots,L$ convolutional layers with $K^{\ell}$ kernels in the $\ell$-th layer, where each kernel is a two-dimensional matrix $r^{\ell}_{x}\times r^{\ell}_{z}$ (since the data is observed in two dimensions). The kernels each perform a local convolution of the input every $s=2$ elements (strides) and are summed across the kernels from the $(\ell-1)$-th layer. Strided convolutions combine downsampling and feature extraction in a single operation, reducing computational cost compared to separate convolution and pooling layers (Springenberg & Uhl 2019). As such, the $\ell$-th convolutional layer with $K^{\ell}$ kernels and stride $s$ maps yields $K^{\ell}$ feature maps, each one of dimensions $n_{x}^{\ell}\times n_{z}^{\ell}$, i.e., a three-dimensional matrix $n_{x}^{\ell}\times n_{z}^{\ell}\times K^{\ell}$, where $n_{x}^{\ell}=\lfloor n_{x}^{\ell-1}/s\rfloor$ and $n_{z}^{\ell}=\lfloor n_{z}^{\ell-1}/s\rfloor$. In the $\ell$-th layer, denote by $Z^{\ell}_{i,j,f}$ the $(i,j)$ entry of the $f$-th feature map, by $w^{\ell}_{a,b,\tilde{f},f}$ the $(a,b)$ entry of the $f$-th kernel for the $\tilde{f}-$th channel of the previous layer, $\tilde{f}\in\{1,\ldots,K^{\ell-1}\}$, and by $w^{\ell}_{0,f}$ the bias term. The spatial encoder performs the following:
$Z^{\ell}_{i,j,f}=g\left(\sum_{\tilde{f}=1}^{K^{\ell-1}}\sum_{a=1}^{r_{x}^{\ell}}\sum_{b=1}^{r_{z}^{\ell}}Z^{\ell-1}_{si+a,sj+b;\tilde{f}}\cdot w^{\ell}_{a,b;\tilde{f},f}+w^{\ell}_{0,f}\right).$ (2)
If we denote the weights in the $\ell$-th layer of the spatial encoder by $\bm{W}^{\ell}=\left\{\left(w^{\ell}_{a,b,\tilde{f},f},w^{\ell}_{0,f}\right):a=1,\ldots,r_{x}^{\ell},~b=1,\ldots,r_{z}^{\ell},~\tilde{f}=1,\ldots,K^{\ell-1},~f=1,\ldots,K^{\ell}\right\}$ and $\bm{Z}^{\ell}=\{Z^{\ell}_{i,j,f}:i=1,\ldots,n^{\ell}_{x},~j=1,\ldots,n^{\ell}_{z},~f=1,\ldots,K^{\ell}\}$, then, the spatial encoder $\mathcal{SE}(\cdot\mid\bm{W}_{\text{en}})$ can be written in matrix form
$\bm{Z}^{\ell}=g\left(\bm{Z}^{\ell-1}\ast\bm{W}^{\ell}\right),~\ell=1,\ldots,L,$ (3)
[…]
$\left|\bm{W}_{\text{en}}\right|=\sum_{\ell=1}^{L}\left(r_{x}^{\ell}\cdot r_{z}^{\ell}\cdot K^{\ell-1}+1\right)\cdot K^{\ell}.$ (4)
The spatial decoder consists of $L+1$ deconvolutional layers, which are similar to convolutional layers but increase the number of dimensions of the feature maps. In the $\ell$-th deconvolutional layer, denote the number of kernels by $\dot{K}^{\ell}$ and the size of each kernel by $\dot{r}_{x}^{\ell}\times\dot{r}_{z}^{\ell}$. In the first $L$ layers of the decoder, the number of kernels in the $\ell$-th layer is the same as the number of kernels in the $(L-\ell-1)$-th layer of the encoder, so the spatial encoder and spatial decoder “mirror” each other, where $(\dot{n}_{x}^{\ell},\dot{n}_{z}^{\ell},\dot{K}^{\ell})=\left(n_{x}^{L-\ell-1},n_{z}^{L-\ell-1},K^{L-\ell-1}\right)$. Since the $L$-th layer of the decoder yields $K^{1}$ feature maps (i.e., the number of feature maps in the first layer of the encoder), the spatial decoder includes one final (output) layer with $\dot{K}^{L+1}=d$ kernels, so that the output matches the dimensions of the data, $n_{x}\times n_{z}\times d$. The deconvolution operation is similar to (2) except it increases the dimensions of the feature maps. Indeed, the $\ell$-th deconvolutional layer with $\dot{K}^{\ell}$ kernels and stride equal to $s$ yields a feature map of dimensions $\dot{n}_{x}^{\ell}\times\dot{n}_{z}^{\ell}\times\dot{K}^{\ell}$, where $\dot{n}_{x}^{\ell}=s\dot{n}_{x}^{\ell-1}$ and $\dot{n}_{z}^{\ell}=s\dot{n}_{z}^{\ell-1}$. The input to the spatial decoder $\mathcal{SD}(\cdot\mid\bm{W}_{\text{de}})$ is the $L$-th layer’s output in the spatial encoder, $\bm{Z}^{L}=\bm{\tilde{Y}}$. In the $\ell$-th layer, if denote by $\dot{Z}^{\ell}_{i,j,f}$ the $(i,j)$ entry of the $f$-th feature map, by $\dot{w}^{\ell}_{a,b,\tilde{f},f}$ the $(a,b)$ entry of the $f$-th kernel for the $\tilde{f}-$th channel of the previous layer, $\tilde{f}\in\{1,\ldots,\dot{K}^{\ell-1}\}$, and by $\dot{w}^{\ell}_{0,f}$ the bias term. The spatial decoder performs the following:
$\dot{Z}^{\ell}_{i,j,f}=g\left(\sum_{\tilde{f}=1}^{\dot{K}^{\ell-1}}\sum_{a=1}^{\dot{r}_{x}^{\ell}}\sum_{b=1}^{\dot{r}_{z}^{\ell}}\dot{Z}^{\ell-1}_{\lfloor\frac{i}{s}\rfloor+a,\lfloor\frac{j}{s}\rfloor+b;\tilde{f}}\cdot\dot{w}^{\ell}_{a,b,\tilde{f},f}+\dot{w}^{\ell}_{0,f}\right).$
If we denote the weights in the $\ell$-th layer of the spatial decoder by $\dot{\bm{W}}^{\ell}=\left\{\left(\dot{w}^{\ell}_{a,b,\tilde{f},f},\dot{w}^{\ell}_{0,f}\right):a=1,\ldots,\dot{r}_{x}^{\ell},~b=1,\ldots,\dot{r}_{z}^{\ell},\tilde{f}=1,\ldots,K^{\ell-1},~f=1,\ldots,\dot{K}^{\ell}\right\}$ and $\dot{\bm{Z}}^{\ell}=\{\dot{Z}^{\ell}_{i,j,f}:i=1,\ldots,\dot{n}^{\ell}_{x},~j=1,\ldots,\dot{n}^{\ell}_{z},~f=1,\ldots,\dot{K}^{\ell}\}$, then, the spatial decoder $\mathcal{SD}\left(\cdot\mid\bm{W}_{\text{de}}\right)$ can be written in matrix form
$\dot{\bm{Z}}^{\ell}=g\left(\dot{\bm{Z}}^{\ell-1}\dot{\ast}\dot{\bm{W}}^{\ell}\right),~\ell=1,\ldots,L,$ (5a)
$\widehat{\bm{Y}}=g\left(\dot{\bm{Z}}^{L}\dot{\ast}\dot{\bm{W}}^{L+1}\right),$ (5b)
[…]
Figure 1: Architecture of the CAE model. The input $\bm{Y}$ is projected onto reduced space through a series of strided convolutions, (2) in yellow, to $\bm{\tilde{Y}}\in\mathbb{R}^{\tilde{n}_{x}\times\tilde{n}_{z}\times\tilde{d}}$ and subsequently projected back onto the original dimensions, $\mathbb{R}^{n_{x}\times n_{z}\times d}$, through a series of strided deconvolutions, in blue.
3.2 Temporal Model: Recurrent Neural Network
The temporal structure of $\bm{Y}_{t}$ is modeled in reduced space, i.e., we model the sequence $\tilde{\mathcal{Y}}^{+}_{t;a}=\left(\tilde{\bm{Y}}_{t},\tilde{\bm{Y}}_{t+1},\ldots,\tilde{\bm{Y}}_{t+a-1}\right)$, as a function of the previous $b$ observations $\tilde{\mathcal{Y}}^{-}_{t;b}=\left(\tilde{\bm{Y}}_{t-b},\tilde{\bm{Y}}_{t-b+1},\ldots,\tilde{\bm{Y}}_{t-1}\right)$. The temporal model comprises of two components: a context builder, which learns the temporal structure of the (reduced space) input sequence into a context representation, and a sequence generator, which learns how to generate the future sequence given the context representation. This sequence-to-sequence approach fundamentally differs from autoregressive strategies that predict only a single future timestep at each iteration and then feed that prediction back into the model to forecast subsequent steps. The problem is well-documented in highly nonlinear systems and multistep forecasting, where autoregressive methods tend to diverge from physically realistic trajectories (Pascanu et al. 2013, Sutskever et al. 2014, Taieb & Atiya 2016, Guen & Thome 2019). In Figure S11 of the supplementary material we provide evidence with our data.
The sequence-to-sequence formulation implemented in this work generates the entire future sequence $\tilde{\mathcal{Y}}^{+}_{t;a}$ in a single forward pass, conditioned on the context of $\tilde{\mathcal{Y}}^{-}_{t;b}$. As such, this approach minimizes temporal inconsistencies due to long-range iterative forecasting. We denote the context builder by $\mathcal{T}_{\text{con}}(\cdot\mid\bm{W}_{\text{con}})$ and the sequence generator by $\mathcal{T}_{\text{gen}}(\cdot\mid\bm{W}_{\text{gen}})$. The modeling approach is similar to that of a language translation model, wherein $\mathcal{T}_{\text{con}}$ learns the “meaning” of $\tilde{\mathcal{Y}}^{-}_{t;b}$ and $\mathcal{T}_{\text{gen}}$ “translates” it into $\tilde{\mathcal{Y}}^{+}_{t;a}$. The temporal model is therefore the composition of the sequence generator and the context builder,
$\mathcal{T}\left(\cdot\mid\bm{W}^{\text{time}}\right)=\mathcal{T}_{\text{gen}}\left(\mathcal{T}_{\text{con}}\left(\cdot\mid\bm{W}_{\text{con}}\right)\mid\bm{W}_{\text{gen}}\right).$
Both the context builder and sequence generator are a ConvLSTM model, an RNN that is able to capture short- and long-term temporal dependencies while also accounting for the spatial structure of the data which, in this case, is a sequence of $b$ three-dimensional matrices, each of dimensions $\tilde{n}_{x}\times\tilde{n}_{z}\times\tilde{d}$ (Hochreiter & Schmidhuber 1997, Shi et al. 2015). At time step $t$, a ConvLSTM model takes as input a (reduced space) observation $\tilde{\bm{Y}}_{t}$ along with the hidden and cell state matrices $\left(\bm{H}_{t-1},\bm{C}_{t-1}\right)$, which carry temporal and memory information from previous time steps, and outputs their respective value for the next time step, formally,
$\bm{I}_{t}$ $=\sigma\left(\bm{W}^{k}_{y,i}\ast\tilde{\bm{Y}}_{t}+\bm{W}^{k}_{h,i}\ast\bm{H}_{t-1}+\bm{W}^{k}_{c,i}\circ\bm{C}_{t-1}+\bm{W}^{k}_{0,i}\right),$ (6a)
$\bm{F}_{t}$ $=\sigma\left(\bm{W}^{k}_{y,f}\ast\tilde{\bm{Y}}_{t}+\bm{W}^{k}_{h,f}\ast\bm{H}_{t-1}+\bm{W}^{k}_{c,f}\circ\bm{C}_{t-1}+\bm{W}^{k}_{0,f}\right),$ (6b)
$\bm{C}_{t}$ $=\bm{F}_{t}\circ\bm{C}_{t-1}+\bm{I}_{t}\circ f\left(\bm{W}^{k}_{y,c}\ast\tilde{\bm{Y}}_{t}+\bm{W}^{k}_{h,c}\ast\bm{H}_{t-1}+\bm{W}^{k}_{0,c}\right),$ (6c)
$\bm{O}_{t}$ $=\sigma\left(\bm{W}^{k}_{y,o}\ast\tilde{\bm{Y}}_{t}+\bm{W}^{k}_{h,o}\ast\bm{H}_{t-1}+\bm{W}^{k}_{c,o}\circ\bm{C}_{t}+\bm{W}^{k}_{0,o}\right),$ (6d)
$\tilde{\bm{H}}_{t}$ $=\bm{O}_{t}\circ f(\bm{C}_{t}).$ (6e)
In (6), $\sigma(\cdot)$ denotes the sigmoid function, $\circ$ denotes the Hadamard product, and $f(\cdot)$ denotes the hyperbolic tangent function. Additionally, $\bm{I}_{t}$ is the input gate activation matrix at time $t$, $\bm{O}_{t}$ is the output gate activation matrix, and $\ast$” denotes the convolution operation (2), in this case with stride $s=1$, kernel size $\tilde{r}_{x}\times\tilde{r}_{z}$, and number of kernels $\tilde{K}$. The parameters to learn are $\bm{W}_{k}=\left\{\bm{W}^{k}_{x,m}:x\in\{y,h,c,0\},~m\in\{i,f,c,o\}\right\},~k\in\{\text{con},\text{gen}\}$.
The temporal structure of the input sequence $\tilde{\mathcal{Y}}^{-}_{t;b}=\left(\tilde{\bm{Y}}_{t-b},\tilde{\bm{Y}}_{t-b+1},\ldots,\tilde{\bm{Y}}_{t-1}\right)$ is learned and stored in context matrices $\left(\bm{H}_{t-1},\bm{C}_{t-1}\right)$, which capture the short- and long-term dynamics of $\tilde{\mathcal{Y}}^{-}_{t;b}$, respectively. As such, the context matrices are initialized $\bm{H}_{t-b-1}=\bm{C}_{t-b-1}=\bm{0}$ and $\mathcal{T}_{\text{con}}$ is defined as in (6) and operates recursively over the input sequence, as described by the following recurrent relationship,
$\left(\tilde{\bm{Y}}_{t-j+1},\bm{H}_{t-j},\bm{C}_{t-j}\right)=\mathcal{T}_{\text{con}}\left(\tilde{\bm{Y}}_{t-j},\bm{H}_{t-j-1},\bm{C}_{t-j-1}\mid\bm{W}_{\text{con}}\right),~j=b,b-1,\ldots,2,1.$ (7)
Importantly, only the final hidden and cell states, $\left(\bm{H}_{t-1},\bm{C}_{t-1}\right)$, are retained and passed to the sequence generator and the output of each step in the context builder, $\tilde{\bm{Y}}_{t-j+1}$, is discarded. This reflects the context builder’s role to summarize the past into context. A visual representation of the context builder is shown in Figure S2 of the supplementary material.
The sequence generator $\mathcal{T}_{\text{gen}}$ learns how to “translate” the context matrices $\left(\bm{H}_{t-1},\bm{C}_{t-1}\right)$ into the (reduced space) output sequence $\tilde{\mathcal{Y}}^{+}_{t;a}=\left(\tilde{\bm{Y}}_{t},\tilde{\bm{Y}}_{t+1},\ldots,\tilde{\bm{Y}}_{t+a-1}\right)$. Therefore, the inputs to the sequence generator are $\left(\tilde{\bm{Y}}_{t-1},\bm{H}_{t-1},\bm{C}_{t-1}\right)$, the last observation of the input sequence and the context matrices. The sequence generator is also implemented as a ConvLSTM and operates recursively as defined in (6), formally,
$\left(\tilde{\bm{Y}}_{t+j},\bm{H}_{t+j},\bm{C}_{t+j}\right)=\mathcal{T}_{\text{gen}}\left(\tilde{\bm{Y}}_{t+j-1},\bm{H}_{t+j-1},\bm{C}_{t+j-1}\mid\bm{W}_{\text{gen}}\right),~j=0,1,\ldots,a-1.$ (8)
Importantly, unlike in (7), the output at each step in (8) is fed back into the sequence generator and used as input for the next prediction. This recursive mechanism allows the model to generate multi-step forecasts in a single forward pass while maintaining temporal consistency. A key advantage of this approach is that it avoids the error accumulation and discontinuities inherent in traditional iterative forecasting methods, where entire predicted sequences are repeatedly fed back into the model. Instead, the ConvLSTM’s hidden and cell states propagate sequential dependencies internally, enabling stable long-horizon predictions. A visual representation of the sequence generator is shown in Figure S2 the supplementary material.
3.3 Spatiotemporal Model: Convolutional Recurrent Neural Network
The final output sequence is obtained by reconstructing the output of the temporal model, $\tilde{\mathcal{Y}}^{+}_{t;a}$, back onto the original dimensions using the spatial decoder (5). Given input and output sequences, $\mathcal{Y}^{-}_{t;b}=\left(\bm{Y}_{t-b},\bm{Y}_{t-b+1},\ldots,\bm{Y}_{t-1}\right)$ and $\mathcal{Y}^{+}_{t;a}=\left(\bm{Y}_{t},\bm{Y}_{t+1},\ldots,\bm{Y}_{t+a-1}\right)$, respectively, our state-space model (1) is then
$\tilde{\bm{Y}}_{t-i}=\mathcal{SE}\left(\bm{Y}_{t-i}\mid\bm{W}_{\text{en}}\right),\quad i=b,b-1,\ldots,2,1,$
$\left(\cdot,\bm{H}_{t-1},\bm{C}_{t-1}\right)=\mathcal{T}_{\text{con}}\left(\tilde{\mathcal{Y}}^{-}_{t;b},\bm{H}_{t-b-1},\bm{C}_{t-b-1}\ \middle|\ \bm{W}_{\text{con}}\right),$
$\left(\bm{\tilde{Y}}_{t+j},\bm{H}_{t+j},\bm{C}_{t+j}\right)=\mathcal{T}_{\text{gen}}\left(\bm{\tilde{Y}}_{t+j-1},\bm{H}_{t+j-1},\bm{C}_{t+j-1}\ \middle|\ \bm{W}_{\text{gen}}\right),\quad j=0,1,\ldots,a-1$
$\widehat{\mathcal{Y}}^{+}_{t;a}=\mathcal{SD}\left(\tilde{\mathcal{Y}}^{+}_{t;a}\mid\bm{W}_{\text{de}}\right).$
We hereto refer to the spatiotemporal model as $\mathcal{ST}(\cdot\mid\bm{W})$, $\bm{W}=(\bm{W}^{\text{space}},\bm{W}^{\text{time}})$. As a neural network with a convolutional (spatial) and a recurrent (temporal) component, this model is a CRNN, which may be summarized as
[…]
4 Inference
Inference is performed in two stages, wherein we first estimate the spatial parameters $\bm{W}^{\text{space}}$ and then the temporal parameters $\bm{W}^{\text{time}}$ conditional on $\widehat{\bm{W}}^{\text{space}}$. We adopt this strategy since the CAE and RNN serve two distinct (and purposely decoupled) functions. Indeed, the CAE can focus on reconstructing spatial structure, while the RNN can focus on modeling the temporal evolution in the reduced space. While joint training may offer the benefit of task-specific representations optimized end-to-end for forecasting performance, we found in practice that decoupled training yields accurate reconstructions and forecasts, while being computationally tractable. Section 4.1 describes the training and testing set, Section 4.2 describes the chosen model architecture and hyperparameters, Section 4.3 discusses the inferential process for the CAE, and Section 4.4 presents the training strategy for the temporal portion of the CRNN conditional on the estimated CAE. Inference is performed in Python 3.12.0 and Tensorflow 2.13. The PI-CRNN model was trained using 8 CPU cores for general processing and a dedicated NVIDIA A10 GPU for accelerated computation.
4.1 Training and Testing set
The first n=2,000n=2{,}000 (of ouf T=2,500T=2{,}500) observations were used for model training, while the remaining Tn=500T-n=500 were used for testing. For the CAE, the training and testing set are $\mathcal{D}_{\text{train}}^{\text{CAE}}=\left\{\bm{Y}_{1},\bm{Y}_{2},\ldots,\bm{Y}_{n}\right\}$ and $\mathcal{D}_{\text{test}}^{\text{CAE}}=\left\{\bm{Y}_{n+1},\bm{Y}_{n+2},\ldots,\bm{Y}_{T}\right\}$, respectively. For the RNN, since it is a sequence-to-sequence model, its input and output is a sequence of observations, so its training set is constructed by dividing $\mathcal{D}_{\text{train}}^{\text{CAE}}$ into $\tilde{n}$ input and output sequences, $\mathcal{D}_{\text{train}}^{\text{RNN}}=\left\{\left(\mathcal{Y}^{-}_{t;b},\mathcal{Y}^{+}_{t;a}\right):t=b,b+o,b+2o,\ldots,n-a\right\}\subset\mathcal{D}_{\text{train}}^{\text{CAE}}$ and $\mathcal{D}_{\text{test}}^{\text{RNN}}=\left\{\left(\mathcal{Y}^{-}_{t+n;b},\mathcal{Y}^{+}_{t+n;a}\right):t=b,b+o,b+2o,\ldots,n-a\right\}\subset\mathcal{D}_{\text{test}}^{\text{CAE}}$, where $o$ is the offset between sequences. The number of training sequences is therefore $\tilde{n}=\frac{n-a-b}{o}+1$, and we set $o=8$, the smallest $o$ that could fit in memory. Note that all the sequences in $\mathcal{D}_{\text{train}}^{\text{RNN}}$ are subsets of $\mathcal{D}_{\text{train}}^{\text{CAE}}$. Having first trained the CAE, we can represent each input sequence in $\mathcal{D}_{\text{train}}^{\text{RNN}}$ in the reduced space and obtain $\tilde{\mathcal{Y}}^{-}_{t;b}=\left(\tilde{\bm{Y}}_{t-b},\tilde{\bm{Y}}_{t-b+1},\ldots,\tilde{\bm{Y}}_{t-1}\right)$, where $t=b,b+o,b+2o,\ldots,n-a$, so that the training and testing sets for the temporal model are $\tilde{\mathcal{D}}_{\text{train}}^{\text{RNN}}=\left\{\left(\tilde{\mathcal{Y}}^{-}_{t;b},\tilde{\mathcal{Y}}^{+}_{t;a}\right):t=b,b+o,b+2o,\ldots,n-a\right\}$ and $\tilde{\mathcal{D}}_{\text{test}}^{\text{RNN}}=\left\{\left(\tilde{\mathcal{Y}}^{-}_{t+n;b},\tilde{\mathcal{Y}}^{+}_{t+n;a}\right):t=b,b+o,b+2o,\ldots,n-a\right\}$.
4.2 Hyperparameters
In the CAE, we assume that the spatial encoder and decoder each have $L=4$ layers, and the number of kernels, which are all $r_{x}^{\ell}=r_{z}^{\ell}=3$, are $K_{1}=K_{2}=K_{3}=K_{4}=64$, so the reduced space representation for any observation $\bm{Y}_{t}\in\mathbb{R}^{n_{x}\times n_{z}\times d}$ is $\tilde{\bm{Y}}_{t}\in\mathbb{R}^{16\times 16\times 64}$, $n_{x}=n_{z}=256$, resulting in a dimension reduction of 93.75%. The activation function $g$, in (3) and (5), is the leaky rectified linear unit, a modified version of the rectified linear unit that allows some negative values, in every layer except for the last, where $g$ is the hyperbolic tangent function function (Goodfellow et al. 2016). In the RNN, the context builder, $\mathcal{T}_{\text{con}}$, and sequence generator, $\mathcal{T}_{\text{gen}}$, each have $\tilde{K}=128$ kernels, where each kernel is $\tilde{r}_{x}=\tilde{r}_{z}=1$, i.e., the temporal evolution is modeled “pixel-wise” in the reduced space. For the lengths of the input and output sequences, $b$ and $a$, respectively, we leveraged the turnover time parameter $\tau_{c}$, defined as the time required for a fluid particle to complete one full convection cycle, which for this flow is $\tau_{c}=5.4s$ (or 54 observations). We selected $b=80$ and $a=60$ to ensure that both the input and output windows span at least one turnover time. This provides the model with enough temporal context to capture the dominant dynamical structures and evolve them meaningfully into the future. So the temporal model (in reduced space) takes an input sequence of length $b=80$ time steps, $\tilde{\mathcal{Y}}^{-}_{t;b}$, and outputs the next $a=60$ (reduced) observations, $\tilde{\mathcal{Y}}^{+}_{t;a}$. Further details on the formal definition of $\tau_{c}$ are available in Section S.2 of the supplementary material.
4.3 Convolutional Autoencoder Inference
In CAE training, each observation is treated independently, since the objective is to compress any one observation into a space of lower dimensions. We find $\widehat{\bm{W}}^{\text{space}}$ by minimizing the mean squared error with respect to the data,
$\widehat{\bm{W}}^{\text{space}}=\operatorname*{arg\,min}_{\bm{W}^{\text{space}}}\text{MSE}_{\text{space}}\left(\bm{W}^{\text{space}}\right)=\operatorname*{arg\,min}_{\bm{W}^{\text{space}}}\left\{\sum_{i}\lambda^{\text{space}}_{i}\mathcal{L}^{\text{space}}_{i}\left(\bm{W}^{\text{space}}\right)\right\},$ (9)
where $\bm{\lambda}^{\text{space}}=\left\{\lambda^{\text{space}}_{i}:i=u,w,p,\theta\right\}$ are weights balancing the four terms’ respective contributions. The terms $\mathcal{L}^{\text{space}}_{i}$ are the mean square error with respect to each of the four variables. Denote by $\bm{Y}_{t}\left(\bm{W}^{\text{space}}\right)=\left(u_{t}\left(\bm{W}^{\text{space}}\right),w_{t}\left(\bm{W}^{\text{space}}\right),p_{t}\left(\bm{W}^{\text{space}}\right),\theta_{t}\left(\bm{W}^{\text{space}}\right)\right)$ the reconstruction at $t$ as a function of $\bm{W}^{\text{space}}=\left(\bm{W}_{\text{en}},\bm{W}_{\text{de}}\right)$, where $\bm{Y}_{t}\left(\bm{W}^{\text{space}}\right)=\mathcal{SD}\left(\mathcal{SE}\left(\bm{Y}_{t}\mid\bm{W}_{\text{en}}\right)\mid\bm{W}_{\text{de}}\right)$. The loss term for $\mathcal{L}^{\text{space}}_{u}$ is then
$\mathcal{L}^{\text{space}}_{u}\left(\bm{W}^{\text{space}}\right)=\frac{1}{n}\sum_{t=1}^{n}\left[u_{t}-u_{t}\left(\bm{W}^{\text{space}}\right)\right]^{2},$
and $\mathcal{L}^{\text{space}}_{w}$, $\mathcal{L}^{\text{space}}_{p}$, and $\mathcal{L}^{\text{space}}_{\theta}$ are similar.
The $\bm{\lambda}^{\text{space}}$ hyperparameters balance the model training and are not chosen via cross-validation. They are instead implemented via a dynamic method that automatically balances the four terms in (9) and lessens the risk that the optimization ends up at a local minimum (Wang et al. 2023). This ensures that the model does not disproportionately focus on optimizing one output at the expense of others, thereby promoting more balanced training. The approach follows established practices in multi-task learning, where adaptive loss weighting has been shown to improve convergence and final model performance (Wang et al. 2020). Technical details are available in Section S.4 of the supplementary material.
4.4 Recurrent Neural Network Physics-Informed Inference
We estimate the temporal parameters $\bm{W}^{\text{time}}$ conditional on $\widehat{\bm{W}}^{\text{space}}$, the estimated CAE parameters. We also incorporate the RBC governing equations in the inferential process, making it a physics-informed model, so henceforth we refer to it as “PI-CRNN”. Formally, we find $\widehat{\bm{W}}^{\text{time}}$ by minimizing a penalized target function:
$\widehat{\bm{W}}^{\text{time}}$ $=\operatorname*{arg\,min}_{\bm{W}^{\text{time}}}\mathcal{L}\left(\bm{W}^{\text{time}}\right)$
$=\operatorname*{arg\,min}_{\bm{W}^{\text{time}}}\left\{\sum_{i}\lambda^{\text{Data}}_{i}\mathcal{L}^{\text{Data}}_{i}\left(\bm{W}^{\text{time}}\right)+\sum_{j}\lambda_{j}^{\text{PDE}}\mathcal{L}^{\text{PDE}}_{j}(\bm{W}^{\text{time}})\right\},$ (10)
where $i\in\{u,w,p,\theta\}$ and $j\in\{\text{Mass},~\text{Mom-}u,~\text{Mom-}w,~\text{Energy}\}$. The term $\mathcal{L}^{\text{Data}}_{i}$ represents the loss term with respect to the observed data, e.g.,
$\mathcal{L}^{\text{Data}}_{u}\left(\bm{W}^{\text{time}}\right)=\frac{1}{a\cdot\tilde{n}}\sum_{t=1}^{\tilde{n}}\sum_{r=1}^{a}\left(u_{t;r}-u_{t;r}\left(\bm{W}^{\text{time}}\right)\right)^{2},$
where $u_{t;r}$ is the $r$-th element of the output sequence (out of $\tilde{n}$ training sequences) starting at $t$ (in this case, for the horizontal component of velocity) and $u_{t;r}\left(\bm{W}^{\text{time}}\right)$ its prediction as a function of $\bm{W}^{\text{time}}$. The terms $\mathcal{L}^{\text{PDE}}_{j}$ instead represent the loss terms with respect to the governing PDEs, e.g.,
$\mathcal{L}^{\text{PDE}}_{\text{Mass}}\left(\bm{W}\right)$ $=\frac{1}{a\cdot\tilde{n}}\sum_{t=1}^{\tilde{n}}\sum_{r=1}^{a}\left(\frac{\partial u_{t;r}\left(\bm{W}^{\text{time}}\right)}{\partial x}+\frac{\partial w_{t;r}\left(\bm{W}^{\text{time}}\right)}{\partial z}\right)^{2},$
and the vector $\bm{\lambda}^{\text{time}}=\left(\lambda^{\text{Data}}_{i},\lambda^{\text{PDE}}_{j}\right),~i\in\{u,w,p,\theta\},~j\in\{\text{Mass},~\text{Mom-}u,~\text{Mom-}w,~\text{Energy}\}$, contains weights balancing the eight terms’ respective contributions to the objective function $\mathcal{L}(\bm{W}^{\text{time}})$. The derivatives in the PDE penalty are computed using a central difference method, which we detail in the supplementary material. The dynamic loss balancing for $\bm{\lambda}^{\text{time}}$ in (10) is similar to (9) and detailed in Section S.4 of the supplementary material.
We hereto denote the forecast at time $t$ by $\widehat{\bm{Y}}_{t}=\left(\widehat{u}_{t},\widehat{w}_{t},\widehat{p}_{t},\widehat{\theta}_{t}\right)$ to simplify the notation. After training the PI-CRNN model, we predict new realizations of RBC using input sequences in the testing set $\mathcal{D}^{\text{RNN}}_{\text{test}}$ as an input and forecasting over a time horizon of $R$ time steps. The procedure to generate forecasts is outlined in Algorithm 1.
Algorithm 1 Forecasting sequences of arbitrary length
[…]
4:$t=0$
5:for $t<R$ do $\triangleright$ $R$ is the forecast horizon
6:  $\bm{\widehat{\tilde{Y}}}_{t},\bm{H}_{t},\bm{C}_{t}=\mathcal{T}_{\text{gen}}\left(\bm{\widehat{\tilde{Y}}}_{t-1},\bm{H}_{t-1},\bm{C}_{t-1}\ \middle|\ \widehat{\bm{W}}_{\text{gen}}\right)$ $\triangleright$  Reduced space prediction and context matrices
7:  $\widehat{\bm{Y}}_{t}=\mathcal{SD}\left(\bm{\widehat{\tilde{Y}}}_{t}\mid\widehat{\bm{W}}_{\text{de}}\right)$ $\triangleright$ Forecast (decompression to the original spatial scale)
8:  $t=t+1$
9:end for
[…]
5 Uncertainty Quantification
The high sensitivity of RBC with respect to the initial condition makes predicting its specific dynamics exceptionally challenging. Infinitesimally small perturbations, e.g., in thermal fluctuations or boundary layer instabilities, create diverging dynamics, thereby limiting the reliability of uncertainty quantification (UQ) over long forecasts (Leutbecher & Palmer 2008, Ahlers et al. 2009, Lorenz 1963). For this reason, we limit UQ to the first two turnover times of the forecast (corresponding to $T_{\text{uq}}=108$ time steps), beyond which the rapid divergence of trajectories renders probabilistic forecasts impractical.
We quantify the uncertainty using a bootstrap-based calibration method, which relies on conformal prediction intervals (Gammerman et al. 1998, Bonas & Castruccio 2023, Bonas et al. 2024). Conformal prediction intervals are constructed by estimating residual uncertainty using quantiles, thereby not relying on the Gaussianity assumption. Since the model was trained up to $n$, we generate an ensemble of $E=50$ forecasts over two turnover times ($T_{\text{uq}}$ total time steps) given input sequences $\mathcal{Y}^{-}_{t;b}$, $n\leq t\leq T-T_{\text{uq}}$. So, the $T_{\text{uq}}$ predicted observations are all within the testing set. Denote the $j$-th element of the long-range forecast starting at $e$ by $\widehat{\bm{Y}}_{t_{e}+j}$, $e=1,\ldots,E$. The $E$ forecasts are then grouped according to their position within the estimated sequence, $\bm{f}_{1}=\left(\widehat{\bm{Y}}_{t_{1}+1},\widehat{\bm{Y}}_{t_{2}+1},\ldots,\widehat{\bm{Y}}_{t_{E}+1}\right)$, $\bm{f}_{2}=\left(\widehat{\bm{Y}}_{t_{1}+2},\widehat{\bm{Y}}_{t_{2}+2},\ldots,\widehat{\bm{Y}}_{t_{E}+2}\right)$, etc… So $\bm{f}_{1}$ contains all the one-step ahead forecasts, $\bm{f}_{2}$ contains all the two-step ahead forecasts, and so on. Then, we compute the absolute value of the residuals for each $t$-step ahead, $\bm{r}_{t}=\left(\bm{R}_{t_{1}+t},\bm{R}_{t_{1}+t},\ldots,\bm{R}_{t_{E}+t}\right)$, where $\bm{R}_{t_{e}+t}=\left|\bm{Y}_{t_{e}+t}-\widehat{\bm{Y}}_{t_{e}+t}\right|,~\bm{R}_{t_{e}+t}=\left(\bm{R}_{u,t_{e}+t},\bm{R}_{w,t_{e}+t},\bm{R}_{p,t_{e}+t},\bm{R}_{\theta,t_{e}+t}\right),~t=1,\ldots,T_{\text{uq}},~e=1,\ldots,E$. The samples $\bm{r}_{t}$ are used to estimate the $\alpha$-quantile $\bm{q}_{j,t}$, where $j=u,w,p,\theta$, formally,
$\bm{\hat{q}}_{u,t}=\inf\left\{\bm{r}_{u,t}:\frac{1}{E}\sum_{e=1}^{E}\mathbb{I}\left(\bm{R}_{u,t_{e}+t}\leq\bm{r}_{u,t}\right)\geq 1-\alpha\right\},~\bm{\hat{q}}_{u,t}\in\mathbb{R}^{n_{x}\times n_{z}},$
and similarly for $\bm{\hat{q}}_{w,t}$, $\bm{\hat{q}}_{p,t}$, and $\bm{\hat{q}}_{\theta,t}$, $\hat{\bm{q}}_{t}=\left(\bm{\hat{q}}_{u,t},\bm{\hat{q}}_{w,t},\bm{\hat{q}}_{p,t},\bm{\hat{q}}_{\theta,t}\right)$. This conformal prediction framework thus enables the construction of valid uncertainty intervals that naturally expand over time, reflecting increasing forecast uncertainty in RNN predictions. By grouping the $E$ forecasts based on their lead time, the method allows residuals to be computed separately for each time step in the forecast horizon, capturing the model’s historical error behavior at that specific lead time. Lastly, we perform an additional calibration step, for each variable, by multiplying the interval by a factor $\bm{\zeta}=\left(\zeta_{u},\zeta_{w},\zeta_{p},\zeta_{\theta}\right)$, found via cross validation (Bonas et al. 2024). So, $(1-\alpha)$% prediction intervals can be obtained by construction as $P\left(\widehat{\bm{Y}}_{t}-\bm{\zeta}\bm{\hat{q}}_{t}\leq\bm{Y}_{t}\leq\widehat{\bm{Y}}_{t}+\bm{\zeta}\bm{\hat{q}}_{t}\right)\approx(1-\alpha)\%$.
6 Results
We now present the results of our approach and first compare its spatial performance against competing statistical approaches and then the temporal one. In regard to the dimension-reduction (and subsequent reconstruction) of the data, we compare the CAE with principal component analysis (PCA), independent component analysis (ICA), fixed rank Kriging (FRK), and proper orthogonal decomposition (POD). We compare the forecast obtained using our PI-CRNN with a non-physics-informed CRNN, an echo state network (ESN), and an autoregressive integrated moving average model (ARIMA). We also assess the ability of PI-CRNN to forecast well beyond the training forecast horizon.
We assess our model’s performance according to its two stated objectives: efficient reduced space representation and long-range forecasting of RBC. For spatial reconstruction, we analyze the CAE’s ability to represent high-dimensional spatial data into a reduced space and subsequently reconstruct it back onto the original dimensions. The metrics we employ are the mean squared error and the structural similarity index measure (SSIM, Wang et al. (2004)) between the reconstruction and ground truth over the testing set. The SSIM was originally conceived to quantify the similarity between images by comparing visual fidelity with respect to the original data. For long-range forecasting, we assess adherence to the governing PDE as well as physical properties of RBC that should be preserved. We focus on 1) the Nusselt number: a measure of how efficiently heat is transferred through a fluid by convection as compared to conduction; 2) the probability density functions (PDFs) of the velocity component and temperature; 3) dissipation rate of thermal fluctuations: loss of temperature variance due to heat diffusion and conduction in the fluid, respectively (Emran & Schumacher 2012). The forecast we generate is of 7 turnovers $\tau_{c}$, or 37.8 seconds, i.e., $R=378$, which is considerably longer (hence more challenging) than previous attempts to forecast similar turbulent flows (Mohan et al. 2019, Mo et al. 2019, Tang et al. 2020). A thorough explanation on the reliability of the forecast as a function of time is available in Section S.6 of the supplementary material. Technical details on each metric of assessment are also available in Section S.6 of the supplementary material. Due to RBC’s extreme sensitivity to initial conditions, even minute discrepancies, e.g., numerical perturbations, can cause forecasts to diverge rapidly over time. This inherent unpredictability makes uncertainty quantification essential. Rather than relying on point predictions alone, it is crucial to capture the range of plausible future states the system may evolve into. Therefore, in our work we also aim at assessing the quality and reliability of the predictive uncertainty provided by our PI-CRNN.
The following sub-sections are organized as follows. In Section 6.1 we analyze the spatial portion of our model, while in Section 6.2 we assess the PI-CRNN forecast and its adherence to the underlying physics. Finally, in Section 6.3 we discuss the long-range forecasting ability of PI-CRNN against existing numerical methods in terms of physical accuracy and computational cost.
6.1 Spatial Analysis (Dimension Reduction)
The reconstruction performance of our CAE presents an alternative to established reduced order modeling techniques in fluid mechanics, which seek to ease the computational burden in numerical simulations without sacrificing key physical properties of the system of interest. The complexity of a dynamical system such as RBC is often reduced with basic decomposition methods. Such approaches predicate the decomposition of a random vector field $\bm{Y}(x,t)$ into a finite sum, e.g., $\bm{Y}(x,t)=\sum_{i=1}^{M}a_{i}(t)\phi_{i}(x)$, where $\phi_{i}(x)$ are the spatial basis functions and $a_{i}(t)$ are temporal coefficients, such that the the approximation improves as $M\rightarrow\infty$. Common choices of $\phi_{i}$ include Fourier series, Legendre polynomials, and Chebyshev polynomials (Chatterjee 2000). Our CAE model accomplishes the same objective in nonlinear fashion. This approach makes the physical interpretation of $\bm{\tilde{Y}}_{t}$ more challenging, but provides increased flexibility granted by the nonlinear activation function $g$ in the convolution step (3) and deconvolution step (5).
In Table 1 we report the $\text{MSE}_{\text{space}}$ and SSIM, along with the interquartile range (IQR) for the CAE, PCA, ICA, FRK and POD. For PCA and ICA we chose the highest possible number of principal components $n$ and used them to reconstruct the TnT-n validation observations. For a fair comparison with CAE, in the case of FRK we interpolated in space (independently for each realization of RBC) using every fourth grid point, since the CAE reduces the original dimensions by a factor of 16. For POD, we chose the maximum possible number of modes. Details on PCA, ICA, FRK and POD are available in Section S.6.1 of the supplementary material. In Figure 2 we show one validation observation $\bm{Y}_{t}$ and our reconstructed estimate $\widehat{\bm{Y}}_{t}$.
The results clearly highlight that other approaches are suboptimal in capturing the nonlinear nature of this data, as the CAE yields a validation MSE of $\mathcal{O}(10^{-6})$, FRK of $\mathcal{O}(10^{-4})$, and PCA and ICA of $\mathcal{O}(10^{-3})$. FRK delivers the second best reconstruction MSE, which is over 10 times worse than that of the CAE, most likely driven by the spatial nonstationarity of RBC in the vertical direction. The suboptimal performance of PCA and ICA, both linear methods, can be attributed to the highly nonlinear nature of the data, though PCA has been shown to be effective in RBC simulations with lower Rayleigh numbers, i.e., less turbulence (Pandey et al. 2022). The results in terms of SSIM follow the same pattern of $\text{MSE}_{\text{space}}$. The relatively high SSIM for FRK, PCA, and ICA can be attributed to a fundamental difference in how it evaluates reconstruction quality compared to $\text{MSE}_{\text{space}}$. SSIM assesses perceptual similarity, making it more sensitive to the preservation of global patterns and spatial coherence rather than small-scale errors. Our result is significant as it reveals how effective the CAE is in extracting the spatial features of RBC while at the same time reducing the dimensions of the data by a factor of 16. As such, reconstructions whose MSE is $\mathcal{O}(10^{-6})$ imply that the CAE retains the key features in the reduced space representation, as is apparent from Figure 2.
Dimension-Reduction
Method $\text{MSE}_{\text{space}}$ (IQR) $\times 10^{-3}$ SSIM (IQR)
CAE $\bm{6.62\times 10^{-3}~\left(8.70\times 10^{-2}\right)}$ $\bm{0.99~\left(2.36\times 10^{-4}\right)}$
FRK $1.65\times 10^{-1}~\left(1.86\times 10^{-2}\right)$ 0.98 $\left(2.86\times 10^{-3}\right)$
POD $1.85~\left(1.32\right)$ 0.97 $\left(2.81\times 10^{-2}\right)$
PCA $2.34~\left(1.52\right)$ 0.96 $\left(3.39\times 10^{-2}\right)$
ICA $4.17~\left(3.51\right)$ 0.96 $\left(4.81\times 10^{-2}\right)$
Table 1: Comparison of dimension-reduction methods in terms of MSE and SSIM on the validation set. In parenthesis we report the IQR across the spatial domain.
Figure 2 illustrates the performance of the CAE on a representative example from the validation set at t=2,125t=2{,}125. The top row displays the original data fields for the four variables ($u$, $w$, $p$, $\theta$, from left to right), while the bottom row shows the corresponding reconstructions obtained from the CAE after reducing the data dimensionality by 93.75%. Despite this significant compression, the CAE successfully preserves the main spatial structures and fine-scale patterns. In particular, the reconstruction captures the large-scale flow structures, sharp gradients, and coherent features present in the original data. This provides strong visual evidence for the model’s ability to faithfully represent the essential dynamics of RBC. The high fidelity of the reconstructions further supports the quantitative results, as reflected by the low $\text{MSE}_{\text{space}}$ and the high SSIM from Table 1.
Data Refer to caption
Reconstruction Refer to caption
Figure 2: CAE results for one observation in the validation set, t=2,125t=2{,}125, showing the data (top row) and the reconstruction (bottom row) after reducing the dimensions of the data by 93.75%. The $\text{MSE}_{\text{space}}$ and SSIM for this observation are $6.12\times 10^{-6}$ and $0.99$, respectively.
6.2 Temporal Analysis
We compare our forecast of $\tau_{c}=7$ turnover times with the ESN, CRNN, and ARIMA. The CRNN model is identical to PI-CRNN, but is not informed with the governing PDE, so the loss function we minimize is similar to (10) but without the “PDE” terms. The ESN is a type of recurrent neural network with a similar structure as a LSTM network, but with considerably less computational demand dictated by the sparseness of its parameter matrices (McDermott & Wikle 2019). The main advantage of ESNs is the absence of backpropagation through time (which is the computational bottleneck of LSTM and ConvLSTM), since only the output weights are trained. Since we perform physics-informed inference, we hereby denote it by “PI-ESN”. The forecast is generated by recursively applying the ESN $R$ times. Lastly, the ARIMA model is one of the most popular approaches in time series analysis thanks to its simplicity and ease of implementation. Similar to the ESN, we generate a forecast of length $R$ by recursively applying ARIMA. Technical details on PI-ESN, CRNN, and ARIMA are available in Section S.6.2 of the supplementary material.
The results in Table 2 report the comparison on forecast performance against the aforementioned methods, showing (I) $\text{MSE}_{\text{RBC}}$, the mean squared error with respect to the governing PDE, (II) the mean squared error with respect to the estimate of the Nusselt number (Nu), (III) the Kullback-Leibler (KL) divergence between the true and estimated PDFs of $\widehat{u}_{t}$, $\widehat{w}_{t}$, and $\widehat{\theta}_{t}$, and (IV) the mean squared error with respect to the dissipation rate of thermal fluctuations $\varepsilon^{\theta}$. In each case, the IQR (computed across forecasts with $E=50$ different starting sequences) is reported in parenthesis.
Our PI-CRNN yields the forecast that most closely resembles a time series of RBC that is physically-consistent, with a $\text{MSE}_{\text{RBC}}$ of $0.01~\left(3.12\times 10^{-3}\right)$, which is approximately 97% smaller than the second best competing approach, CRNN. While PI-ESN attains the second-lowest $\text{MSE}_{\text{RBC}}$, its PDFs deviate considerably from ground truth, indicating systematic biases in its representation of coherent structures. CRNN ranks third in terms of $\text{MSE}_{\text{RBC}}$, yet it better preserves physical realism, producing more accurate distributions and rare-event statistics, making it a stronger overall model than PI-ESN. Similarly, the key physical attributes of the forecast, represented here by the Nusselt number, probability density functions of the velocity and temperature, and the dissipation rate of thermal fluctuations, are best preserved in PI-CRNN. The error with respect to a perfectly physical forecast, captured by $\text{MSE}_{\text{RBC}}$, increases over time, as small-scale discrepancies propagate through time. Nonetheless, overall statistical adherence is mostly preserved, as indicated by a small mean squared error with respect to the true Nusselt number. The dissipation is a much stricter benchmark, since it depends on accurately capturing small-scale gradients, and it presents a more difficult challenge for PI-CRNN to capture. The CRNN model yields the second best results, with a $\text{MSE}_{\text{RBC}}$ of 0.74 $\left(9.39\times 10^{-2}\right)$, primarily thanks to the same sequence-to-sequence approach implemented in PI-CRNN, though performance degrades quickly due to the absence of direct embedding of the physics in the inferential process. Indeed, Nu and the PDFs, both of which require a long (physically-consistent) forecast in order to obtain reasonable estimates, are markedly worse. The PI-ESN model yields results that are comparable to CRNN, but it suboptimally captures the distributional characteristics of RBC. This is likely due to fact that ESNs’ reservoir weights are in fact random and untrained, limiting their ability to control information flow over long horizons. The worst performing model is (expectedly) ARIMA, which is unable to capture the highly chaotic, nonlinear nature of RBC. In Section S.6 of the supplementary material we report a comprehensive comparison of these three physical metrics. Figures S5-S7 show the Nusselt number, PDFs, and dissipation, respectively. Figure S10 of the supplementary material show additional visual evidence for the superior performance of PI-CRNN.
(I) (II) (III) (IV)
[…]
PI-CRNN $\bm{0.01~\left(3.12\times 10^{-3}\right)}$ $\bm{22.11~(40.45)}$ $\bm{5.05~(2.91)}$ $\bm{3.55\times 10^{-3}~\left(1.08\times 10^{-3}\right)}$
CRNN 0.74 $\left(9.39\times 10^{-2}\right)$ 2,544.28 (1,277.46) 10.86 (8.92) $4.89\times 10^{-3}~\left(1.99\times 10^{-3}\right)$
PI-ESN 0.08 $\left(5.50\times 10^{-2}\right)$ 31.14 (406.02) 2613.94 (375.12) $1.02\times 10^{-2}~\left(5.81\times 10^{-3}\right)$
ARIMA 7.34 $\left(1.39\times 10^{-1}\right)$ 710.75 (7,214.42) 3755.95 (823.93) $2.94\times 10^{-1}~\left(7.92\times 10^{-3}\right)$
Table 2: Comparison of temporal methods in terms of physical metrics: (I) is the mean squared error with respect to the governing equations, (II) is the mean squared error with respect to the true Nusselt number, (III) is the Kullback-Leibler divergence between the forecast’s PDFs and those of the data, and (IV) is the mean squared error with respect to the dissipation rate of thermal fluctuations.
Figure 3 presents a sequence of snapshots from the temperature forecast $\widehat{\theta}_{t}$ over the course of the first two turnover times, corresponding to approximately 11 seconds of physical time. Each panel shows the predicted temperature field at different times, highlighting the evolution of the thermal structures. Two fully developed plumes can be clearly observed rising from the bottom boundary and interacting with the surrounding flow. PI-CRNN is able to capture the key dynamical features of plume formation, growth, and lateral motion over time. Although the mean squared error with respect to the true data increases gradually with forecast lead time, indicating some accumulation of small deviations, the residuals of the governing equations remain relatively constant, especially over the first two turnover times (see Figure S10 of the supplementary material). This suggests that while the forecast slowly diverges from the realized RBC for $t>n$, the model consistently preserves the underlying physical constraints. An animation showing the full temporal evolution is available in the online code associated with the manuscript.
Figure 3: Snapshots of the temperature forecast over the first turnover time ($\approx$ 5.5s), where the spatiotemporal dynamics remain strongly dependent on the input sequence. An animation is available in the online code associated with the manuscript.
6.3 PI-CRNN as a Surrogate
We now assess the PI-CRNN efficacy as a statistical surrogate to produce physically consistent, long-range forecasts. As summarized in Table 3, the model demonstrates strong generative capabilities, as we are able to produce $\tau_{c}=7$ turnover times worth of physical realizations of future RBC states in under 20 seconds. By comparison, the runtime required for a direct numerical simulation (DNS, the numerical approach used to generate the training data) is more than five times larger. In DNS, partial differential equations are integrated in time by evaluating spatial and temporal derivatives at each time step, meaning that each new realization must be computed sequentially from previous ones. By contrast, our PI-CRNN decouples the learning of flow dynamics from explicit numerical integration by embedding the structure of the governing RBC equations into the parameter estimates through the physics-informed training process outlined in Section 4.4. PI-CRNN does not require expensive physical solvers, enabling its use as an “off-the-shelf” surrogate capable of producing statistically faithful forecasts of RBC.
Method Cost Per Step Entire Simulation
Direct simulation 0.20 s 75.6 s
PI-CRNN 0.04 s 14.5 s
Table 3: Comparison of turbulence generation methods in terms of computational demand. We show the per-step computational time and the total for $R=378$ time steps.
It must be noted that PI-CRNN does not yield results with exact adherence to the governing PDE at every time step. Small deviations of the surrogate forecast from the original simulation accumulate over time, but they are mitigated by penalizing against non-physical solutions, leading to the retention of key statistical properties of RBC, e.g., the Nusselt number, as discussed in Section 6.2. In fact, even with seemingly identical initial conditions, ensemble members of a DNS are expected to diverge from one another in an instantaneous sense (though they will still adhere to the governing equations). This divergence, which arises from the chaotic nature of fluid flows, helps to motivate our understanding of the key statistical features of RBC, rather than instantaneous features associated with individual ensemble members. By contrast, a purely data-driven approach such as CRNN (defined in Section 6.2), does not preserve any physical properties of RBC and deviations from the governing PDE grow significantly over time. Additional details and results for CRNN are provided in the supplementary material, where we show visual evidence for the superior physical performance of PI-CRNN over CRNN.
7 Conclusion
In this work we present a spatiotemporal, physics-informed convolutional recurrent neural network model, PI-CRNN, which acts as a surrogate to generate long-range forecasts of RBC convection—a buoyancy-driven flow that manifests between surfaces exhibiting a temperature gradient. Simulating RBC convection by solving the underlying PDEs is time consuming and presents scalability challenges. Our proposed model is considerably faster and offers a scalable and efficient alternative for long-range forecasting of RBC.
Our approach compresses the spatial representation of RBC using a convolutional autoencoder, and models its temporal evolution using a recurrent neural network within a sequence-to-sequence framework inspired by natural language processing. This approach fundamentally differs from classical strategies that predict a single future timestep and iteratively feed the extended sequence back into the temporal model. In contrast, the sequence-to-sequence formulation generates the entire future sequence in a single forward pass, conditioned on the context of the input sequence, thereby minimizing the risk of compounding errors. The approach is also physics-informed, as model parameters are estimated via penalized optimization with respect to the governing equations of RBC.
PI-CRNN’s features, namely its convolutional structure, recurrent memory, and physics-based loss, are generalizable to other nonlinear systems in thermodynamics, climate, and geophysical flows. While the current implementation focuses on fixed physical parameters, our framework is readily extensible to different parameter regimes. This is particularly important for capturing the wide range of turbulent flow behaviors that arise under varying levels of convection. Mesoscale processes emerging from nonlinear interactions between small-scale dynamics and larger structures such as thermal plumes could also benefit from such an approach.
Future work will generalize PI-CRNN by incorporating additional spatiotemporally-aware neural network architectures suitable for high-dimensional data, such as vision transformers (Dosovitskiy et al. 2021). Unlike convolutional neural networks that rely on local spatial relationships, vision transformers use self-attention mechanisms to capture global spatial dependencies across the entire domain. This capability allows them to efficiently capture long-range interactions between distant regions, a key feature in thermodynamic systems, where large-scale structures such as plumes and vortices exert system-wide influence.
Code and Data Availability
The code for this work can be found in the following GitHub repository: github.com/papersblindreview/PI_CRNN_RBC_2025. The numerical solver for the RBC used in this work can be freely downloaded from git.uwaterloo.ca/SPINS/SPINS_main.
Acknowledgments
Stefano Castruccio was partially funded by NSF grant OAC-2347239 and the Notre Dame Provost Strategic Framework Grants. David Richter was partially funded by NSF grant AGS-2227012. Andrew Grace was partially funded by the Engineering Research Council of Canada fund number PDF - 587364 - 2024.